# Non-iterative Optimization of Trajectory and Radio Resource for Aerial Network

### v1

@@ Abstract
We address a joint trajectory planning, user association, resource allocation, and power control problem to maximize proportional fairness in the aerial IoT network, considering practical end-to-end quality-of-service (QoS) and communication schedules. Though the problem is rather ancient, apart from the fact that the previous approaches have never considered user- and time-specific QoS, we point out a prevalent mistake in coordinate optimization approaches adopted by the majority of the literature. Coordinate optimization approaches, which repetitively optimize radio resources for a fixed trajectory and vice versa, generally converge to local optima when all variables are differentiable. However, these methods often stagnate at a non-stationary point, significantly degrading the network utility in mixed-integer problems such as joint trajectory and radio resource optimization. We detour this problem by converting the formulated problem into the Markov decision process (MDP). Exploiting the beneficial characteristics of the MDP, we design a non-iterative framework that cooperatively optimizes trajectory and radio resources without initial trajectory choice. The proposed framework can incorporate various trajectory planning algorithms such as the genetic algorithm, tree search, and reinforcement learning. Extensive comparisons with diverse baselines verify that the proposed framework significantly outperforms the state-of-the-art method, nearly achieving the global optimum. Our implementation code is available at https://github.com/hslyu/dbspf.
@@ I Introduction
These use cases of aerial networks align closely with current research focus on internet-of-things (IoT) networks, as documented in the 3GPP white papers [3, 4]. One of the most critical agendas for realizing aerial IoT networks involves communication scheduling to compensate for the constrained memory and energy capacity [5, 6].
Most research detours the variable entanglement by adopting coordinate optimization approaches which reciprocally optimize TP variables for fixed radio resource variables and vice versa [7]. In the optimization-theoric viewpoint, these approaches are well-known to converge to local optima in differentiable problems, but may halt at the non-stationary point [^1] A point is stationary if changes in any direction make utility degradation (e.g. local maxima). A point is non-stationary if the point is not stationary. when optimizing mixed-integer problems, even if the objective is convex [8].
We have found an interesting, but catastrophic interaction between trajectory-planning (TP) and radio resource management (RRM), characterized by “non-optimal initialization can trap the network variables in the non-stationary point.” Figure 1 illustrates the curse of initialization that could deteriorate the network utility unless the initial variable choice is optimal, which is rarely the case. The discovery suggests that a surprising number of research works may inadvertently optimize the utility of aerial networks in a possibly incorrect manner [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24].
We ask ourselves “How can we optimize aerial networks while avoiding the curse of initialization?” and find the answer in the non-iterative approach. As in Fig. 1, we first convert radio resources as a function of trajectory. Then, we find the optimal trajectory by integer programming [^2] We assume a discretized trajectory for the non-convex, non-differentiable problem. For a continuously differentiable trajectory, convex relaxations and optimizations can be applied.. The optimal RRM is, in turn, determined by the trajectory, resulting in the optimal solution.
@@
  • Resolve the curse of initialization: This research is, to the authors’ knowledge, the first to highlight that iterative optimization might be inherently unhelpful in aerial networks. We address the problem through the MDP reformulation, which provides a straightforward, powerful non-iterative framework by separately optimizing TP and RRM problems. Figure 2 briefly compares the result of the iterative optimization and the proposed method. In Sec. V, we show that the proposed method outperforms various comparison schemes and validate the robustness of the proposed method.
  • Propose a novel temporal decoupling method: We suggest a novel method to separate the PF problem into unit-time sub-problems. The PF entangles all controllable variables in the logarithm over the entire time steps. Then, applying well-known RRM techniques, specialized to optimize unit-time network snapshots, becomes challenging. We devise multiple mathematical tricks to convert the logarithm of sum-rate (PF formulation) into a summation of logarithms. Standing on the temporal decoupling, we show that the proposed RRM scheme tightly achieves the global optimum.
@@
Figure 2: Motivated illustration of our contribution. Graphs (a) and (b) represent iterative optimization (GA-ITER), and graph (c) represents the proposed method (GA-TP). The two methods share the same environment configurations and algorithms except for the optimization order. While GA-ITER optimizes radio resources for a given initial trajectory, GA-TP adaptively optimizes radio resources according to the changing trajectory. Note that the initial trajectory in (a) restricts the updated trajectory in (b), resulting in a large difference between the result trajectories in (b) and (c). Detail descriptions are provided in the Sec. V.
@@ II System Model
Figure 3: Illustration of the system model at time $t=20$. The dark gray marks represent users whose request periods have already expired. The green marks with dashed links represent users getting serviced at the current time. The orange marks represent users whose request periods do not expire, but that are not currently being served. Users who have not expired may not be served if their request period does not start. The UAV-BS selects service users by jointly optimizing TP, UA, RA, and PC.
@@ II-A Scenario Description
When the $t$-th time slot begins, the UAV-BS must be located on a certain grid point of the flight map. We denote $\mathbf{q}^{(t)}$ as the position of the UAV-BS at the beginning of the $t$-th time slot, where $\mathbf{q}^{(t)}$ is defined as
@@ 2 Output $\mathbf{A}^{(t)}$, $\mathbf{B}^{(t)}$
3 $\mathbf{A}_{\text{current}}^{(t)}\leftarrow\{\}$.
@@
5       $\mathcal{I}_{\text{feasible}}\leftarrow\{i\in\mathcal{I}|\sum_{i\in\mathbf{A}_{\text{current}+i}^{(t)}}r_{i}/e_{i}^{(t)}\leq B\}$
6       $\mathbf{A}_{\rm{next}}^{(t)}\leftarrow$ The best UA set among $\mathbf{A}_{\rm{current}}^{(t)}$ and $\mathbf{A}_{\rm{current}+i}^{(t)},~{}i\in\mathcal{I}_{\rm{feasible}}$.
7
@@
Now, the objective (21a) can be maximized by finding the optimal $\mathbf{A}^{(t)}$ as the optimal $\mathbf{B}^{(t)}$ can be obtained for arbitrary $\mathbf{A}^{(t)}$, We adopt an incremental algorithm that finds a local-optimal solution (Alg. 1). The key idea of the algorithm is to sequentially find a better UA combination than the current UA combination $\mathbf{A}_{\text{current}}^{(t)}$.
Let $\mathbf{A}_{\text{current}}^{(t)}$ be a UA set and $\mathbf{A}_{\text{current}+i}^{(t)}$ contains one more user $i\notin\mathbf{A}_{\text{current}}^{(t)}$, denoted as
$\mathbf{A}_{\text{current}+i}^{(t)}=\mathbf{A}_{\text{current}}^{(t)}\cup\{i\}.$ (23)
If $\mathbf{A}_{\text{current}+i}^{(t)}$ satisfies constraint (21b) and (21c), $\mathbf{A}_{\text{current}+i}^{(t)}$ could be a maximizer of Problem (21). Two constraints (21b) and (21c) can be simplified, so we can define a set of feasible additive users as
$\mathcal{I}_{\text{feasible}}=\big{\{}i\in\mathcal{I}|\Sigma_{i\in\mathbf{A}_{\text{current}+i}^{(t)}}r_{i}/e_{i}^{(t)}\leq B\big{\}}.$ (24)
Then, we can define a new set $\mathbf{A}_{\text{next}}^{(t)}$ that makes the objective (21a) be the greatest among the set $\mathbf{A}_{\text{current}}^{(t)}$ and $\mathbf{A}_{\text{current}+i}^{(t)}$ for $i\in\mathcal{I}_{\text{feasible}}$. If $\mathbf{A}_{\text{next}}^{(t)}$ is equal to $\mathbf{A}_{\text{current}}^{(t)}$, set $\mathbf{A}_{\text{next}}^{(t)}$ is a solution. Otherwise, $\mathbf{A}_{\text{current}}^{(t)}$ is updated to $\mathbf{A}_{\text{next}}^{(t)}$ and the above procedure is repeated until finding the solution.
By iteratively updating the UA set from the empty set $\mathbf{A}_{\text{current}}^{(t)}=\{\}$, the proposed algorithm can find the local-optimal UA and RA variables.
@@ 4 Function $f(\mathbf{q}^{(t)})$:
7             Update $\mathbf{B}^{(t)}$ by using Alg. 3.
8             Update $\mathbf{P}^{(t)}$ by using (61).
@@ IV-A4 Computational Complexity
The complexity of Alg. 1 is $\mathcal{O}(I^{2})$ as the maximum number of comparisons (Line 6 in Alg. 1) is $\frac{I(I+1)}{2}$. The RA scheme in Alg. 3 utilizes gradient descent with the complexity of $\mathcal{O}(\frac{I}{\epsilon})$ and the closed-form solution (61) takes $\mathcal{O}(1)$. Then, the complexity of the Alg. 2 is $\mathcal{O}(\frac{I^{2}}{\epsilon^{2}})$ since the iterations from Line 6 to Line 9 in Alg. 2 takes $\mathcal{O}(\frac{1}{\epsilon})$.
@@ IV-B Decision-Making Algorithms for Trajectory-Planning
The complexity of the three TP schemes combined with the RRM scheme is computed by multiplying the complexity of the TP scheme and Alg. 2. When $m$ generations and $n$ populations are given, the complexity of GA is $\mathcal{O}(mn\frac{I^{2}}{\epsilon^{2}})$. For DFS, if the maximum number of reachable grid points $m$ with search depth $n$ is defined as $m=\max_{\mathbf{q}\in\mathcal{Q}}|S(\mathbf{q})|$, the complexity of DFS is $\mathcal{O}(mn\frac{I^{2}}{\epsilon^{2}})$. The complexity of DQN is $\mathcal{O}(\frac{I^{2}}{\epsilon^{2}})$ as the feed-forward of the network takes $\mathcal{O}(1)$.
@@ VI Conclusion
This paper proposes a separation of control and RRM in the time-critical aerial network. We find a critical drawback of conventional joint iterative optimizations, where the initial trajectory choice fetters the subsequent variable choices. However, separating TP and RRM in the practical scenario is challenging because the user requirements and fairness tightly bind network variables. We address the challenge by decomposing the proposed problem into the sum of serial sub-problems. Then, we cleverly transform the sub-problems into the MDP formulation where the TP and RRM variables are separately optimized. The proposed methods achieve the global optimum obtained by the genetic algorithm, outperforming the state-of-the-art scheme.
The proposed approach focuses on designing a non-differentiable trajectory by integer programming. We underscore that applying the non-iterative approach to build a differentiable trajectory is another important open problem. Then continuous decision-making algorithms such as deep deterministic policy gradient [49] should be adopted.
Optimizing multi-UAV networks through the MDP formulation is another open challenge. A proliferation of research on multi-UAV networks has adopted iterative optimization, and there might be a gap in network utility due to the initial trajectories (or position) choice problem.
We anticipate that the overall network utility could be enhanced by applying the non-iterative philosophy.
@@ Appendix D Optimal Resource Allocation
where $\gamma$ is a learning rate and $o_{i}^{(t)}=\lambda_{1}^{(t)}+\lambda_{2}^{(t)}\rho_{i}^{(t)}-\mu_{i}^{(t)}$. The initial values of the multipliers are configured [^9] We heuristically apply the bisection method to find the best combination that minimizes the convergence time. as $\mu_{i}^{(t)}=0.5$ for all $i\in\mathcal{I}$, $\lambda_{1}^{(t)}=0.5$, and $\lambda_{2}^{(t)}=0.8$. After that, the optimal value of $\mathbf{B}^{(t)}$ can be obtained from

### v2

@@ Abstract
We address a joint trajectory planning, user association, resource allocation, and power control problem to maximize proportional fairness in the aerial IoT network, considering practical end-to-end quality-of-service (QoS) and communication schedules. Though the problem is rather ancient, apart from the fact that the previous approaches have never considered user- and time-specific QoS, we point out a prevalent mistake in coordinate optimization approaches adopted by the majority of the literature. Coordinate optimization approaches, which repetitively optimize radio resources for a fixed trajectory and vice versa, generally converge to local optima when all variables are differentiable. However, these methods often stagnate at a non-stationary point, significantly degrading the network utility in mixed-integer problems such as joint trajectory and radio resource optimization. We detour this problem by converting the formulated problem into the Markov decision process (MDP). Exploiting the beneficial characteristics of the MDP, we design a non-iterative framework that cooperatively optimizes trajectory and radio resources without initial trajectory choice. The proposed framework can incorporate various trajectory-planning algorithms such as the genetic algorithm, tree search, and reinforcement learning. Extensive comparisons with diverse baselines verify that the proposed framework significantly outperforms the state-of-the-art method, nearly achieving the global optimum. Our implementation code is available at https://github.com/hslyu/dbspf.
@@ I Introduction
These use cases of aerial networks are closely aligned with current research focus on internet-of-things (IoT) networks, as documented in the 3GPP white papers [3, 4]. One of the most critical agendas for realizing aerial IoT networks involves communication scheduling to compensate for the constrained memory and energy capacity [5, 6].
Most research detours the variable coupling by adopting coordinate optimization approaches which reciprocally optimize TP variables for fixed radio resource variables and vice versa [7]. In the optimization-theoric viewpoint, these approaches are well-known to converge to local optima in differentiable problems, but may halt at the non-stationary point [^1] A point is stationary if changes in any direction make utility degradation (e.g. local maxima). A point is non-stationary if the point is not stationary. when optimizing mixed-integer problems, even if the objective is convex [8].
We have found an interesting, but catastrophic interaction between trajectory-planning (TP) and radio resource management (RRM), where “non-optimal initialization can trap the network variables in the non-stationary point.” Figure 1 illustrates how the initialization deteriorates the network utility, unless the initial variable choice is optimal. The discovery suggests that a surprising number of research works may inadvertently optimize the utility of aerial networks in a possibly incorrect manner [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24].
@@
We ask ourselves “How can we optimize aerial networks while avoiding the curse of initialization?” and find the answer in the non-iterative approach. As in Fig. 1, we first convert radio resources as a function of trajectory. Then, we find the optimal trajectory by integer programming. [^2] We assume a discretized trajectory for the non-convex, non-differentiable problem. For a continuously differentiable trajectory, convex relaxations and optimizations can be applied. The optimal RRM is, in turn, determined by the trajectory, resulting in the optimal solution.
@@
Figure 2: Motivated illustration of our contribution. Graphs (a) and (b) represent iterative optimization (GA-ITER), and graph (c) represents the proposed non-iterative approach (GA-TP). The two methods share the same environment configurations and algorithms except for the optimization order. While GA-ITER optimizes radio resources for a given initial trajectory, GA-TP adaptively optimizes radio resources according to the changing trajectory. Note that the initialization in (a) limits the changes of the updated trajectory, leading to a significant difference between the resulting trajectories in (b) and (c).
Detail descriptions are provided in the Sec. V.
@@
  • Resolve the curse of initialization: To the authors’ knowledge, This research is the first to highlight that iterative optimization might be inherently unhelpful in aerial networks. We address the problem through the MDP reformulation, which provides a straightforward, powerful non-iterative framework by separately optimizing TP and RRM problems. Figure 2 briefly compares the result of the iterative optimization and the proposed method. In Sec. V, we show that the proposed method outperforms various comparison schemes and validate the robustness of the proposed method.
  • Propose a novel temporal decoupling method: We suggest a novel method to separate the PF problem into unit-time sub-problems. The PF entangles all controllable variables in the logarithm operations over the entire time steps. Then, applying well-known RRM techniques, specialized to optimize unit-time network snapshots, becomes challenging. We devise multiple mathematical tricks to decompose the logarithm of sum-rate (PF formulation) into a summation of logarithms. Standing on the temporal decoupling, we show that the proposed RRM scheme tightly achieves the global optimum.
@@ II System Model
Figure 3: Illustration of the system model at time $t=20$. The dark gray marks represent users whose request periods have already expired. The green marks with dashed links represent users who are being serviced at the current time. The orange marks represent users whose request periods do not expire, but that are not currently being served. Users who have not expired may not be served if their request period does not start. The UAV-BS selects service users by jointly optimizing TP, UA, RA, and PC.
@@ II-A Scenario Description
The UAV-BS is assumed to be located on a certain grid point of the flight map at the beginning of the $t$-th time slot. The discrete trajectory—also known as waypoint-based navigation—has been a de facto standard in a wide range of aerial vehicles, including commercial airplanes, copters, and fixed-wing vehicles [34, 35]. We denote $\mathbf{q}^{(t)}$ as the position of the UAV-BS at the beginning of the $t$-th time slot, where $\mathbf{q}^{(t)}$ is defined as
@@ 2 Output $\mathbf{A}^{(t)}$, $\mathbf{B}^{(t)}$
3 $\mathbf{A}_{\rm{current}}^{(t)}\leftarrow\{\}$.
@@
5       $\mathbf{A}_{\rm{current}}^{(t)}\leftarrow\mathbf{A}_{\rm{next}}^{(t)}$
6       $\mathcal{I}_{\rm{feasible}}\leftarrow\{i\in\mathcal{I}|\sum_{i\in\mathbf{A}_{\mathrm{current}+i}^{(t)}}r_{i}/e_{i}^{(t)}\leq B\}$
7       $\mathbf{A}_{\rm{next}}^{(t)}\leftarrow$ The best UA set among $\mathbf{A}_{\rm{current}}^{(t)}$ and $\mathbf{A}_{\mathrm{current}+i}^{(t)}~{}\text{for}~{}i\in\mathcal{I}_{\rm{feasible}}$.
8
@@ 5       $\mathbf{A}_{\rm{current}}^{(t)}\leftarrow\mathbf{A}_{\rm{next}}^{(t)}$
Now, the objective (21a) can be maximized by finding the optimal $\mathbf{A}^{(t)}$ as the optimal $\mathbf{B}^{(t)}$ can be obtained for arbitrary $\mathbf{A}^{(t)}$, we adopt an incremental algorithm that finds a local-optimal solution (Alg. 1). The key idea of the algorithm is to sequentially find a better UA combination than the current UA combination $\mathbf{A}_{\mathrm{current}}^{(t)}$.
Let $\mathbf{A}_{\mathrm{current}}^{(t)}$ be a UA set and $\mathbf{A}_{\mathrm{current}+i}^{(t)}$ contains one more user not in $\mathbf{A}_{\mathrm{current}}^{(t)}$, denoted as
$\mathbf{A}_{\mathrm{current}+i}^{(t)}=\mathbf{A}_{\mathrm{current}}^{(t)}\cup\{i\}.$ (23)
If $\mathbf{A}_{\mathrm{current}+i}^{(t)}$ satisfies constraint (21b) and (21c), $\mathbf{A}_{\text{current}+i}^{(t)}$ could be a maximizer of Problem (21). Two constraints (21b) and (21c) can be simplified, so we can define a set of feasible additive users as
$\mathcal{I}_{\mathrm{feasible}}=\big{\{}i\in\mathcal{I}|\Sigma_{i\in\mathbf{A}_{\mathrm{current}+i}^{(t)}}r_{i}/e_{i}^{(t)}\leq B\big{\}}.$ (24)
Then, we can define a new set $\mathbf{A}_{\text{next}}^{(t)}$ that makes the objective (21a) be the greatest among the set $\mathbf{A}_{\mathrm{current}}^{(t)}$ and $\mathbf{A}_{\mathrm{current}+i}^{(t)}$ for $i\in\mathcal{I}_{\mathrm{feasible}}$. If $\mathbf{A}_{\mathrm{next}}^{(t)}$ is equal to $\mathbf{A}_{\mathrm{current}}^{(t)}$, set $\mathbf{A}_{\mathrm{next}}^{(t)}$ is a solution. Otherwise, $\mathbf{A}_{\mathrm{current}}^{(t)}$ is updated to $\mathbf{A}_{\mathrm{next}}^{(t)}$ and the above procedure is repeated until finding the solution.
By iteratively updating the UA set from the empty set $\mathbf{A}_{\mathrm{current}}^{(t)}=\{\}$, the proposed algorithm can find the local-optimal UA and RA variables.
@@ 4 Function $f(\mathbf{q}^{(t)})$:
7             Update $\mathbf{B}^{(t)}$ by using Alg. 3 in Appendix D.
8             Update $\mathbf{P}^{(t)}$ by using (69) in Appendix E.
@@ IV-A4 Computational Complexity
The complexity of Alg. 1 is $\mathcal{O}(I^{2})$ as the maximum number of comparisons (Line 6 in Alg. 1) is $\frac{I(I+1)}{2}$. The RA scheme in Alg. 3 in Appendix D utilizes gradient descent with the complexity of $\mathcal{O}(\frac{I}{\epsilon})$ and the closed-form solution (69) in Appendix E takes $\mathcal{O}(1)$. Then, the complexity of the Alg. 2 is $\mathcal{O}(\frac{I^{2}}{\epsilon^{2}})$ since the iterations from Line 6 to Line 9 in Alg. 2 takes $\mathcal{O}(\frac{1}{\epsilon})$.
@@ IV-B Decision-Making Algorithms for Trajectory-Planning
The complexity of the three TP schemes combined with the RRM scheme is computed by multiplying the complexity of the TP scheme and Alg. 2. When $m$ generations and $n$ populations are given, the complexity of GA is $\mathcal{O}(mn\frac{I^{2}}{\epsilon^{2}})$. For DFS, if the maximum number of reachable grid points $m$ with search depth $n$ is defined as $m=\max_{\mathbf{q}\in\mathcal{Q}}|S(\mathbf{q})|$, the complexity of DFS is $\mathcal{O}(mn\frac{I^{2}}{\epsilon^{2}})$. The complexity of DQN is $\mathcal{O}(\frac{I^{2}}{\epsilon^{2}})$ as the feed-forward of the network takes $\mathcal{O}(1)$. The key factor affecting computational complexity is the number of possible actions $|s(\mathbf{q})|$ in the MDP, as the TP algorithms are derived from the MDP formulation. In large-scale networks, the complexity of the TP algorithms can be reduced to a manageable level by adjusting the grid size $\Delta Q$ and the time slot length $\Delta T$.
@@ VI Conclusion
This paper proposes a separation of control and RRM in the time-critical aerial network. We find a critical drawback of conventional joint iterative optimizations, where the initial trajectory choice fetters the subsequent variable choices. However, separating TP and RRM in the practical scenario is challenging because the user requirements and fairness tightly bind network variables. We address the challenge by decomposing the proposed problem into the sum of serial sub-problems. Then, we cleverly transform the sub-problems into the MDP formulation where the TP and RRM variables are separately optimized. The proposed methods closely achieve the global optimum obtained by the genetic algorithm, outperforming the state-of-the-art schemes.
The proposed approach focuses on designing a non-differentiable trajectory by integer programming. We underscore that applying the non-iterative approach to build a differentiable trajectory is another important open problem. We attach the preliminary results on continuous network optimization conducted on the proposed environment in Appendix G. Then continuous decision-making algorithms such as deep deterministic policy gradient [51] should be adopted.
Optimizing multi-UAV networks through the MDP formulation is another open challenge. A proliferation of research on multi-UAV networks has adopted iterative optimization, and there might be a gap in network utility due to the curse of the initial choice.
We expect that the network utility of aerial networks can be improved by incorporating the non-iterative approach.
@@ Appendix D Optimal Resource Allocation
[…]
@@ Appendix F DQN Architecture and Training Procedure
(a) Discrete trajectory obtained from GA-TP.
(b) Trajectory interpolated by the cubic spline.
Figure 16: Trajectories optimized with 10 moving users and 10 stationary users.
Figure 17: Proportional fairness and number of served users with various QoS constraints in the dynamic environment.
Appendix G Studies on the Continous Optimizations with Moving Users
We additionally design an continuous TP and RRM to verify how the proposed method operates in the realistic scenario. In this scenario, we additionally apply a uniform velcocity to half of the mobile users, considering future developments where IoT devices, such as autonomous cars, AAM/UAMs, and quadrupedal robots, will incorporate mobility. In the following experiments, we assume 10 users are mobile while the other 10 remain stationary. [^10] The velocity of users is randomly assigned with the speed ranging from $[0,30]$ m/s.
We first design a differentiable trajectory by interpolating the result trajectory obtained from the DFS method as shown in Fig. 16.111111The trajectory is interpolated using cubic splines. Then, we measured the PF in two cases: i) reusing the RRM obtained from the DFS algorithm (continuous trajectory w/o RRM); and ii) performing the RRM every 10 ms (continuous trajectory with RRM).
Figure 18: The proportional fairness per time slot for a single episode.
Fig. 17 indicate the PF and %UE over QoS constratins in the above environmental configurations. There is no significant change in trend compared to Figs. 9 and 10. However, slight decreases—1-2 for PF and 2-4% for %UE—are observed.
By making the TP and RRM continuous, the resulting PF measurement also becomes continuous as in Fig. 18. For “Continuous Trajectory without RRM”, we just apply the RRM of the Discrete Trajectory—$\mathbf{A}^{(t)}$, $\mathbf{B}^{(t)}$, and $\mathbf{P}^{(t)}$ for $t\in\mathcal{T}$—to compute the PF at time $t^{\prime}\in[t,t+1)$. Then, we can observe multiple discrete jumps in PF as the user association discretely changes at every $t$-th time slot. Thus, these discrete jumps disappear as the user association changes more granularly when we re-optimize the RRM every 10 ms for the continuous trajectory in “Continuous Trajectory with RRM”.
Appendix H Wall-Clock Time Benchmarks of the TP Algorithms
TABLE V: Wall-clock time of TP algorithms
Methods GA-TP DFS DQN GA-ITER OFDMA $n=1$ $n=3$ $n=5$ Time (s) 84.73 0.09 1.34 33.34 0.50 22.33 - PF 48.86 46.73 47.97 48.57 47.48 47.85 22.10
Table V shows the wall-clock time benchmarks when a UAV-BS serves users with 2 MHz bandwidth during 20 time slots (which is equivalent to 60 seconds). We omit the elapsed time of OFDMA because we implement the scheme using the genetic algorithm [41], not as described in the original article [11].
The proposed schemes, DFS and DQN, demonstrate significantly faster path design capabilities while achieving higher PF values, compared to GA-ITER (which is an iterative framework). This can be attributed to the fact that the proposed methods avoid “the curse of initialization” without the need for repetitive optimization processes.